OpenStudy (anonymous):

Which numbers should be multiplied to obtain 183^2 - 139^2? 44 and 44 1936 and 1936 44 and 322 33350 and 33628

OpenStudy (loser66):

\[183^2 -139^2=(183+139)(183-139)\] = 322*44